Two busses leave Rochester, at the same time, on the same route, heading to NYC. The local bus averages 36.8 mph and the express bus averages 60.4 mph. How long will it take for the two busses to be 5.9 miles apart?
Found 2 solutions by jorel1380, ccs2011:
Answer by jorel1380(3719)

36.8(n)+5.9=60.4(n)
23.6n=5.9
n=5.9/23.6=.25 hours = 15 minutes..

Answer by ccs2011(207)

Distance = rate*time
Set up equation where express bus has traveled 5.9 miles farther than local bus in same amount of time.
60.4t = 36.8t + 5.9
23.6t = 5.9
t = 5.9/23.6
t = 0.25
In terms of minutes, multiply by 60
0.25*60 = 15
So it will take 15 minutes for the buses to be 5.9 miles apart.